# A Multi-Objective Evaluation Framework for Analyzing Utility-Fairness Trade-Offs in Machine Learning Systems

### v1

@@ III-C Radar Chart: Compact Visualization
We consider the areas in the radar chart as polygons and use the Surveyor’s formula [41] for calculation. Given a polygon $Poly$ with $n$ ordered points as counterclockwise in the Cartesian coordinate system, $v_{i}\in V$, we define $Poly={v_{1},v_{2},..,v_{n}}$ alongside $v_{i}=(x_{i},y_{i})$. As we work in the Polar coordinate system to represent the systems in the radar chart, we define $v_{i}=(r_{i},\Theta_{i})$ where $r_{i}$ is the radius and $\Theta_{i}$ is the angle. In this stage, we need to convert the point in Polar coordinates to the counterpart in Cartesian one by $x_{i}=r_{i}\cos(\Theta_{i})$ and $y_{i}=r_{i}\sin(\Theta_{i})$. After switching to the Cartesian coordinate system, we apply the Surveyor’s formula as shown below:
$\Delta_{Poly}=\frac{1}{2}\left\{\begin{vmatrix}x_{1}&x_{2}\\ y_{1}&y_{2}\end{vmatrix}+..+\begin{vmatrix}x_{n-1}&x_{n}\\ y_{n-1}&y_{n}\end{vmatrix}+\begin{vmatrix}x_{n}&x_{1}\\ y_{n}&y_{1}\end{vmatrix}\right\}$ (10)
where |.||.|| . | is the $2\times 2$ determinant. The calculation of $Area$ is then min-max normalized by $\widehat{\Delta}=\frac{\Delta-min(\Delta)}{max(\Delta)-min(\Delta)}$ to transform the area range to $[0,1]$. Given a pentagon with 5 dimensions as shown in Fig. 9a, the theoretical lower and upper bounds of the area are $0.00$ and $\approx 2.37$, respectively. This is analogous to the concept of the Area Under the Curve (AUC) over Receiver Operator Characteristic (ROC), so an area of $1.00$ is expected to be the best situation for a system. In the illustrative example of Fig. 9a, it can be easily seen from the radar chart that $System1$ outperforms $System2$ in every dimension. This can also be verified by their respective areas of $0.83$ and $0.04$ (Table 9b). We can also clearly observe that $System1$ is closer to the ideal performance than $System2$, relying only on these areas.
(a) Dominance of $System1$ (blue) with respect to $System2$ (red) is clearly visible as it occupies a larger volume of the plot. Normalized MOO indicators or their complement are used to bind axes to the [0,1] scale, improving visual analysis.
@@
System $HV$ $\widehat{ONVG}$ $ONVGR$ $\overline{UD}$ $OS$ $\widehat{\Delta}$
$System1$ 0.93 1.00 0.90 0.85 0.89 0.83
@@
III-D Simulations for Use-cases
@@
The first use-case, UC-1, focuses on black-box testing of $System1$ and $System2$, corresponding to the first scenario (Fig. 1). As both systems have just $1$ non-dominated solution, they have the same $\widehat{ONVG}$ and $ONVGR$ values of $1.00$. In terms of the diversity measurements, it’s not possible then to evaluate both systems as there exist only one non-dominated solution. $System1$ has a higher $HV$ score than $System2$ because its non-dominated solution is further away from the nadir point than the non-dominated solution of the other, resulting in a larger volume for $System1$. Table 10b and Fig. 10a show this comparison quantitatively and qualitatively. The radar chart in Fig. 10a illustrates the performance gap and HV dominance of $System1$ over $System2$. The difference in the area is $\widehat{\Delta}=\widehat{\Delta_{1}}-\widehat{\Delta_{2}}=0.26-0.20=0.06$ and it arises solely from the HV differences between the two systems. In this case, the proposed evaluation framework simplifies the comparison using a single indicator, HV, as the other indicators do not play a role in the model selection decision.
In the second use-case, UC-2, we perform hybrid testing with white ($System1$, Fig. 1) and black box cases ($System2$, Fig. 2). Eight different non-dominated solutions are considered out of $25$, adjusting $\tau$ for $System2$ against a single non-dominated solution for $System1$, as it is not tunable. Table 11b contains the indicator scores for both systems. $System2$ outperforms $System1$ for $\widehat{ONVG}$ as it has more non-dominated solutions but results in a lower $HV$ with a score of $0.09$. Even though $System2$ has distribution and spread indicators, they are not comparable since there is no score generated for $System1$. This can be observed in the radar chart shown in Fig. 11a. We can end up with a decision that $System2$ outperforms $System1$ overall, as seen in the difference of the areas $\widehat{\Delta}=\widehat{\Delta_{2}}-\widehat{\Delta_{1}}=0.12-0.03=0.09$.
@@
(a) Dominance of $System1$ (blue) with respect to $System2$ (red) is clearly visible from the non-overlapped area.
@@
System $HV$ $\widehat{ONVG}$ $ONVGR$ $\overline{UD}$ $OS^{*}$ $\widehat{\Delta}$
$System1$ 0.35 1.00 1.00 - 0 0.26
$System2$ 0.03 1.00 1.00 - 0 0.20
${}^{*}1e-2$
@@
The last use-case, UC-3, considers white-box testing for both $System1$ and $System2$, matching the second scenario (Fig. 2). $10$ and $6$ different non-dominated solutions were considered out of $25$ for $System1$ and $System2$, respectively. As seen in Table 12b, $System1$ outperforms $System2$ in terms of $\widehat{ONVG}$, $ONVGR$, $HV$, $\overline{UD}$, and $OS$. This can also be interpreted through a visual inspection of the radar chart in Fig. 12a. Finally, the areas of both systems confirm this as well, $\widehat{\Delta}=\widehat{\Delta_{1}}-\widehat{\Delta_{2}}=0.25-0.05=0.20$.
We can derive some conclusions about performance indicators based on the observations regarding the use-cases above. Firstly, the black-box case only gives a partial characterization of the assessment since the ML system are not tunable. In this case, we obtain non-discriminative values for ONVG or ONVGR of 1.00 as seen in Fig. 10. Secondly, diversity is another issue when working with a low number of solutions. A system with only one solution does not allow us to evaluate UD and OS as there must be at least two solutions. For instance, UD is evaluated by measuring the distance between two solutions or OS is performed on the distance between the most/least optimal points from the system and the ideal/nadir points from the PF. A single point solution fails to have a score of UD and OS as seen in $System2$ shown in Table 11b. Thirdly, HV may be the first decision point to select the system, which exhibits more PF characteristics, as it covers every aspect of convergence, diversity and capacity. There may be opposite cases between HV and the other indicators as seen in Fig. 12. A good indication of ONVG, ONVGR, UD, and OS does not work as well as evaluating the performance over HV, which has been proven to be a more reliable measurement for PF [37, 39].
(a) Dominance of $System2$ (red) with respect to $System1$ (blue) is visible as it occupies a slightly larger volume of the plot.
@@
System $HV$ $\widehat{ONVG}$ $ONVGR$ $\overline{UD}$ $OS^{*}$ $\widehat{\Delta}$
$System1$ 0.24 0.12 1.00 - 0 0.03
$System2$ 0.09 1.00 0.32 0.61 3 0.12
${}^{*}1e-2$
(b) Quantitative results for (a) as black/white-box.
@@
IV Empirical Validation
In this section, we demonstrate the effectiveness of the proposed evaluation framework empirically. A medical image dataset named the Harvard Glaucoma Fairness (HGF) containing a well-specified binary classification task and metadata with protected demographic attributes was used to showcase the novel evaluation framework in a realistic scenario [38]. The HGF dataset contains cases with a retinal nerve disease called glaucoma as well as samples of healthy patients [42]. Glaucoma is twice as common in Black patients when compared to other races, and more prevalent men [43, 38]. The dataset contains $3300$ 2D retinal nerve fiber layer thickness (RNFLT) maps from three racial groups, namely Asian, Black and White, with a resolution of 200 $\times$ 200 pixels around the optic disk. The glaucoma/non-glaucoma ratio in the dataset is as $53.0\%/47.0\%$. The prevalence of glaucoma in Asians, Blacks, and Whites in the HGF dataset is equibalanced as $33.3\%$, and the gender distribution is as $54.9\%$ and $45.1\%$ for females and males, respectively. The ophthalmologic images from HGF are linked to sensitive attributes for race, gender, age, and ethnicity allowing the evaluation of the utility-fairness trade-off in proposed solutions.
To properly observe the full capabilities of the proposed method, we formulate the utility-fairness trade-off system as a bi-objective optimization problem. More specifically, we implement an ML solution for the demographically fair glaucoma classification using Pareto HyperNetworks (PHNs), which are tailored for Pareto optimality [40, 44]. A PHN is a Neural Network (NN) topology that is able to generate sub-NNs with varying degrees of compromise between multiple objectives. In our setting, each sub-NN corresponds to a different trade-off between utility and fairness, i.e., one sub-model represents a performance with higher utility and lower fairness, another has higher equity sacrificing from the classification accuracy. We construct the PHN topology based on the work from [40] and use ResNet-18 as the backbone alongside 25 different preference vectors that represent the sub-NNs with specific levels of utility and fairness. We optimize the utility and fairness by using weighted loss functions based on Binary Cross-Entropy (BCE) and Equalized Odds (EO), respectively, in a back-propagated manner.
Fig. 13 summarizes the evaluation of both systems considering 3 metrics: classification accuracy (utility) alongside 2 equalized odds differences (EOD), one for race and another for gender. We follow the third analyzed scenario from Section III-A, since both systems can be evaluated through a white-box set up. Fig. 13a (a) highlights the importance of the proposed evaluation framework: even with only two fairness criteria considered besides utility, it is not possible to compare the two systems through visual inspection alone of this 3D plot. On the other hand, the proposed radar chart in Fig. 13b, helps to detect a slight advantage from $System2$ when compared to $System1$. This can be further investigated by performing quantitative analysis in Table 13c. Although there is an inconsistency between HV and the other performance indicators (ONVG, ONVGR and UD), an area of $0.37$ over $0.30$ shows that $System2$ is better-suited for the tuning of utility-fairness trade-offs than $System1$. In this case, it is not possible to differentiate the systems over OS as they both share an spread of $20.3e-$ 2. Finally, it is straightforward to interpret that both systems are far from optimality as they are below an ideal performance of $1.00$.
(a) Dominance of $System1$ (blue) with respect to $System2$ (red) is clearly visible from the non-overlapped area as it occupies a larger volume of the plot.
@@ IV Empirical Validation
System $HV$ $\widehat{ONVG}$ $ONVGR$ $\overline{UD}$ $OS^{*}$ $\widehat{\Delta}$
$System1$ 0.53 1.00 0.40 0.59 9 0.25
$System2$ 0.02 0.60 0.24 0.39 1 0.05
${}^{*}1e-2$
@@
(b) $System2$ (red) slightly dominates $System1$ (blue) with a margin.

### v2

@@ 3.3 Radar Chart: Compact Visualization
We consider the areas in the radar chart as polygons and use the Surveyor’s formula (Braden, 1986) for calculation. Given a polygon Poly with $n$ ordered points as counterclockwise in the Cartesian coordinate system, $v_{i}\in V$, we define $\mathit{Poly}={v_{1},v_{2},\ldots,v_{n}}$ alongside $v_{i}=(x_{i},y_{i})$. As we work in the polar coordinate system to represent the systems in the radar chart, we define $v_{i}=(r_{i},\Theta_{i})$ where $r_{i}$ is the radius and $\Theta_{i}$ is the angle. In this stage, we need to convert the point in polar coordinates to the counterpart in Cartesian one by $x_{i}=r_{i}\cos(\Theta_{i})$ and $y_{i}=r_{i}\sin(\Theta_{i})$. After switching to the Cartesian coordinate system, we apply the Surveyor’s formula as shown below:
$\Delta_{\mathit{Poly}}=\frac{1}{2}\left\{\begin{vmatrix}x_{1}&x_{2}\\ y_{1}&y_{2}\end{vmatrix}+\ldots+\begin{vmatrix}x_{n-1}&x_{n}\\ y_{n-1}&y_{n}\end{vmatrix}+\begin{vmatrix}x_{n}&x_{1}\\ y_{n}&y_{1}\end{vmatrix}\right\}$ (12)
where $|.|$ is the $2\times 2$ determinant. The calculation of Area ($\Delta$) is then min-max normalized by $\widehat{\Delta}=\frac{\Delta-\min(\Delta)}{\max(\Delta)-\min(\Delta)}$ to transform the area range to $[0,1]$. Given a pentagon with 5 dimensions as shown in Fig. 9a, the theoretical lower and upper bounds of the area are $0.00$ and $\approx 2.37$, respectively. This is analogous to the concept of the Area Under the Curve (AUC) over Receiver Operating Characteristic (ROC), where an area of $1.00$ is expected to be the best situation for a system. In the illustrative example of Fig. 9a, it can be easily seen from the radar chart that System1 outperforms System2 in every dimension. This can be verified by their respective areas of $0.84$ and $0.04$ (Table 9b). We can also clearly observe that System1 is closer to the ideal performance than System2, relying only on these areas.
(a) Dominance of System1 (blue) with respect to System2 (orange) is clearly visible as it occupies a larger volume of the plot. Normalized MOO indicators are used to bind axes to the [0,1] scale, improving visual analysis.
@@
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$
$System1$ 0.93 1.00 0.90 0.85 0.89 0.84
@@
3.4 Deduplication of Solutions
In our study, each point on the utility-fairness trade-off curve corresponds to an ML model generated under a specific preference vector. These points form a solution set, $S$, in a multi-objective space, where the dimensions represent utility and fairness metrics. However, multiple preference vectors may cause to build ML models that exhibit nearly identical behavior, and the resulting trade-off system may have (near-)duplicate points in $S$. Such redundant ML models distort performance indicators by artificially increasing density or changing the characteristics of the approximated PF.
To alleviate this issue, we use a deduplication operator, denoted by $\mathit{deduplicate}_{\varepsilon}(.)$, which filters out ML models that lie within $\varepsilon$-neighborhood of each other given a utility-fairness trade-off system. We leverage the DBSCAN clustering algorithm (Schubert et al., 2017) to retain only the representative ML models by eliminating redundant ones. Given $S={s_{1},s_{2},\ldots,s_{M}}\subset\mathcal{R}^{N}$ where $M$ and $N$ refer to the number of solutions and objectives, respectively, applying deduplication operator, $S^{\prime}=\mathit{deduplicate}_{\varepsilon}(S)$, eliminates models in similar performance within $\varepsilon$ formulated as below:
$\lvert s_{i}-s_{j}\rvert>\varepsilon,\quad s_{i},s_{j}\in S^{\prime}$ (13)
We empirically set $\varepsilon=1e-6$ based on experiments conducted on synthetic and real-world datasets.
3.5 A Priori and A Posteriori Analysis
In ML evaluation, it is essential to distinguish between a priori and a posteriori analysis, as each serves a different role in assessing the generalization and stability of the systems in comparison. In the context of demographic fairness, we adapt this concept into our evaluation framework to achieve a similar analysis on utility-fairness trade-off systems compatible with ML assessment.
In our context, a priori analysis refers to identifying Pareto-optimal operating points by using a validation set, prior to observing the final test data. This procedure allows the selection of operating combinations of thresholds and sub-ML models (derived from the trade-off system) that constitute the estimated non-dominated solution set on the validation set. The goal of this analysis is to perform a real deployment scenario in which system parameters must be fixed before test-time evaluation. In contrast, a posteriori analysis relies directly on test-set evaluations and assesses the performance of all sub-ML models without any pre-selection. While a posteriori evaluations provide a complete characterization of the attainable utility-fairness spectrum, they do not represent a real generalization scenario, since the selection step is already considered with test data.
Our framework supports both protocols: a priori evaluation performs a realistic operating mode, whereas a posteriori evaluation provides a full diagnostic understanding of the system’s trade-off characteristics. To have a priori evaluation in practice, we allow users to assess two subsets as validation and test by using the same trained ML model. Based on the validation set, the framework can determine the threshold/sub-ML combinations that define the Pareto-optimal estimate. These selections can then be directly applied to the test set to measure how well the trade-off structure generalizes. In Section 4.3, we provide use cases for both analyses based on real-world medical problems.
3.6 Simulations for Use Cases
@@
The first use case, UC-1, focuses on black-box testing of System1 and System2, corresponding to the first scenario (Fig. 1). As both systems have just $1$ non-dominated solution, they have the same $\widehat{\mathit{ONVG}}$ and ONVGR values of $1.00$ and $0.50$ respectively. In terms of the diversity measurements, it is not possible to evaluate both systems as there exists only $1$ non-dominated solution. System1 has a higher HV score than System2 because its non-dominated solution is farther from the nadir point than the non-dominated solution of the other, resulting in a larger volume. Table 10b and Fig. 10a show this comparison quantitatively and qualitatively. The radar chart in Fig. 10a illustrates the performance gap and HV dominance of System1 over System2. The difference in the area is $\widehat{\Delta}=\widehat{\Delta_{1}}-\widehat{\Delta_{2}}=0.27-0.21=0.06$ and it arises solely from the HV differences between the two systems. In this case, the proposed evaluation framework simplifies the comparison using a single indicator, HV, as the other indicators do not play a role in the model selection decision.
@@
(a) Dominance of System1 (blue) with respect to System2 (orange) is clearly visible from the non-overlapped area.
@@
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$
System1 0.35 1.00 0.50 1.00 0.00 0.27
System2 0.04 1.00 0.50 1.00 0.00 0.21
@@
In the second use case, UC-2, we perform hybrid testing with black- (System1, Fig. 1) and white-box (System2, Fig. 2) cases. $8$ different non-dominated solutions are considered out of $25$, adjusting $\tau$ for System2 against a single non-dominated solution for System1, as it is not tunable. Table 11b contains the indicator scores for both systems. System2 outperforms System1 for $\widehat{\mathit{ONVG}}$ as it has more non-dominated solutions but results in a lower HV with a score of $0.09$. The distribution indicator is not informative as UD is same for both systems, and the spread is better for System2 with an AS of $0.26$ compared to $0.00$ for System1. This can be observed in the radar chart shown in Fig. 11a. We can end up with a decision that System2 outperforms System1 overall, as seen in the difference of the areas $\widehat{\Delta}=\widehat{\Delta_{2}}-\widehat{\Delta_{1}}=0.43-0.12=0.31$.
(a) Dominance of System2 (orange) with respect to System1 (blue) is visible as it occupies a larger volume of the plot.
@@
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$
System1 0.24 0.12 0.50 1.00 0.00 0.12
System2 0.09 1.00 0.89 1.00 0.26 0.43
(b) Quantitative results for (a) as black- and white-box.
@@
The last use case, UC-3, considers white-box testing for both System1 and System2, matching the second scenario (Fig. 2). $10$ and $6$ different non-dominated solutions were considered out of $25$ for System1 and System2, respectively. As seen in Table 12b, System1 outperforms System2 in terms of $\widehat{\mathit{ONVG}}$, ONVGR, HV, and AS with same performance in terms of UD. This can also be interpreted through a visual inspection of the radar chart in Fig. 12a. Finally, the areas of both systems confirm this as well, $\widehat{\Delta}=\widehat{\Delta_{1}}-\widehat{\Delta_{2}}=0.61-0.31=0.30$.
We can derive some conclusions about performance indicators based on the observations regarding the use cases above. Firstly, the black-box case only gives a partial characterization of the assessment since the ML systems are not tunable. In this case, we obtain non-discriminative values for $\widehat{\mathit{ONVG}}$, ONVGR, UD, and AS as seen in Fig. 10. Secondly, diversity is another issue when working with a low number of solutions. A system with only $1$ solution does not allow to evaluate AS as there must be at least two solutions to measure the distance between the extreme points from the system and PF. A single point solution generates an AS score of $0.00$ as seen in System1 shown in Table 11b. Thirdly, HV may be the first decision point to select the system, which exhibits more PF characteristics, as it covers every aspect of convergence, diversity and capacity. There may be opposite cases between HV and the other indicators as seen in Fig. 11. A good indication of $\widehat{\mathit{ONVG}}$, ONVGR, UD, and AS does not work as well as evaluating the performance over HV, which has been proven to be a more reliable measurement for PF (Audet et al., 2021; Jiang et al., 2014).
(a) Dominance of System1 (blue) with respect to System2 (orange) is clearly visible from the non-overlapped area as it occupies a larger volume of the plot.
@@
System $\bm{\mathit{HV}}$ $\bm{\widehat{\mathit{ONVG}}}$ $\bm{\mathit{ONVGR}}$ $\bm{\mathit{UD}}$ $\bm{\mathit{AS}}$ $\bm{\widehat{\Delta}}$
System1 0.54 1.00 0.91 1.00 0.46 0.61
System2 0.02 0.60 0.86 1.00 0.18 0.31
@@
4 Empirical Validation
In this section, we demonstrate the effectiveness of the proposed evaluation framework empirically using three fairness-aware medical imaging datasets.
4.1 Dataset Description
To empirically validate the proposed evaluation framework, we use three medical imaging datasets with demographic attributes: the Harvard Glaucoma Fairness (HGF) dataset (Luo et al., 2024), the Shenzhen Chest X-ray dataset (Jaeger et al., 2014), and the mBRSET retinal dataset (Wu et al., 2025). These datasets were selected because they jointly capture demographic and clinical diversity, and enable the evaluation of utility-fairness trade-offs under both modality-based and population-based conditions.
Harvard Glaucoma Fairness (HGF).
The HGF dataset includes cases with a retinal nerve disease called glaucoma as well as samples of healthy patients. Glaucoma is twice as common in Black patients compared to other races, and more prevalent in men (Khachatryan et al., 2019; Luo et al., 2024). The dataset comprises $3300$ two-dimensional retinal nerve fiber layer thickness (RNFLT) maps from three racial groups, namely Asian, Black, and White, with a resolution of $200\times 200$ pixels. Color intensity represents retinal nerve fiber thickness in micrometers around the optic disc. The glaucoma/non-glaucoma ratio in the dataset is $53.0\%/47.0\%$. The prevalence of glaucoma in Asians, Blacks, and Whites in the HGF dataset is equally balanced at $33.3\%$, and the gender distribution is as $54.9\%$ and $45.1\%$ for females and males, respectively. The ophthalmologic images from HGF are linked to sensitive attributes for race, gender, age, and ethnicity. This dataset provides a clear setting for assessing demographic disparities, as the higher glaucoma prevalence among Black patients contradicts the relative scarcity of such samples in ophthalmology data. This fact creates a fairness challenge that represents real-world imbalance in healthcare AI systems and provides a solid use case for our framework.
Shenzhen Chest X-ray Dataset.
The Shenzhen dataset is a public chest X-ray collection designed for computer-aided screening of pulmonary diseases, particularly tuberculosis (TB). The dataset contains $662$ frontal chest X-rays with a distribution of normal and TB-positive cases as $326$ and $336$, respectively. Image resolutions vary with an approximation of $3000\times 3000$ pixels. The dataset has sensitive attributes for age and gender, providing a complementary use case for evaluating fairness across clinical subgroups with chest radiographs.
mBRSET Retinal Dataset.
The mBRSET dataset is a retinal imaging resource and designed to enable the benchmarking of automated ophthalmologic screening models under different demographic conditions. The dataset consists of $5164$ retinal images from $1291$ subjects, including fundus photography modality. Each sample is annotated for ocular diseases such as diabetic retinopathy and macular edema, alongside demographic attributes such as patient age, gender, and obesity. The image resolution varies in height and width, ranging from $874$ and $951$ to $2304$ to $2984$ pixels, respectively. The dataset exhibits a gender imbalance, with female patients compared to male ones as $65.1\%/34.9\%$. The patient age has an average of $61.4$ years with a standard deviation of $11.6$.
(a) HGF RNFLT map.
(b) Shenzhen X-ray.
(c) mBRSET fundus.
Figure 13: Sample images from the datasets used in the empirical study: (a) RNFLT map from HGF for glaucoma detection; (b) chest X-ray from Shenzhen for pulmonary disease screening; and (c) retinal fundus image from mBRSET for diabetic retinopathy classification.
By jointly employing these three datasets, our empirical evaluation covers complementary fairness perspectives across medical imaging. HGF provides a clear setting for analyzing demographic and disease prevalence imbalances in glaucoma with race and gender labels. The Shenzhen dataset extends this perspective by providing chest X-ray imagery annotated with sensitive attributes such as gender, enabling fairness assessment in a different clinical modality. Finally, mBRSET provides clinical diversity across ocular conditions by having a broader diagnostic context with different retinal diseases. Together, these datasets enable a more comprehensive examination of how the proposed framework evaluates utility-fairness trade-offs across different medical imaging modalities, disease types, and demographic distributions.
4.2 System Definitions and Configuration
To properly demonstrate the capabilities of the proposed evaluation framework, we formulate the utility-fairness trade-off as a bi-objective optimization problem through two distinct systems denoted as System1 and System2, which differ from the systems considered in previous experiments. For the HGF dataset, both systems are developed using Pareto HyperNetworks (PHNs), which are based on hypernetworks (Ha et al., 2016), and are designed to learn Pareto-optimal solutions across multiple objectives (Navon et al., 2020). The PHN formulation allows the generation of sub-neural networks (sub-NNs) that represent different levels of compromise between diagnostic utility and fairness, enabling continuous exploration along the PF. In particular, each sub-NN corresponds to a specific utility-fairness trade-off level, i.e., one sub-model represents a diagnostic performance with higher utility but lower fairness, whereas another has higher equity at the cost of classification accuracy. Each objective combination is represented by a preference vector that encodes the relative weighting between the two objectives. In the HGF setting, all sub-NNs are evaluated directly on the test set, and the full spectrum of utility-fairness outcomes is characterized as a posteriori, without a validation-driven selection of operating points.
Each PHN employs a ResNet-18 backbone as a shared encoder that generates parameters for sub-NNs corresponding to different preference vectors. A total of $25$ preference vectors are uniformly sampled from a Dirichlet distribution with $\alpha=0.2$ to represent distinct utility-fairness trade-offs. The utility objective is defined by Binary Cross-Entropy (BCE) loss, and the fairness objective is modeled using a differentiable relaxation over Equalized Odds (EO) criterion, which captures disparities in true positive and false positive rates (TPR, FPR) across demographic subgroups. The combined objective is jointly optimized through back-propagation to dynamically generate Pareto-optimal solutions that balance diagnostic utility and fairness.
For the mBRSET and Shenzhen datasets, System1 and System2 use two different model architectures. System1 is based on the DenseNet topology (Huang et al., 2017), and this setting follows a training mechanism performed directly under a bi-objective loss combining BCE for utility and EO-based fairness penalties to represent an independent ML model that corresponds to a specific preference. Unlike the joint optimization characteristic of the PHN, here, each utility-fairness trade-off is modeled as a separate ML model. System2 employs a LoRA-enabled ViT-Small model (Hu et al., 2022; Dosovitskiy, 2020), where low-rank adaptation layers are inserted into the self-attention blocks to enable lightweight fairness-aware fine-tuning. For both datasets, the hyperparameters (e.g., learning rate, optimizer, and batch size) follow a fixed configuration within each system. For the mBRSET, PF operating points are selected on a validation set and then evaluated on the test set, following a priori analysis of the utility-fairness trade-offs, whereas for Shenzhen, all operating points are evaluated directly on the validation set, corresponding to an a posteriori analysis.
System1 serves as a baseline fairness configuration across all datasets. It focuses on optimizing fairness with respect to a binary sensitive attribute in each case. For the HGF dataset, this system minimizes EO disparities between male and female patients, targeting gender fairness. For the mBRSET, the system similarly minimizes EO disparities between obese and non-obese patients, capturing obesity-related fairness. For the Shenzhen dataset, the system again focuses on gender fairness by minimizing EO disparities between same subgroups.
System2 extends this baseline to investigate dataset specific demographic variation. In the HGF setting, this system shifts the fairness objective from gender to race by minimizing EO disparities between Asian, Black, and White subgroups. For the mBRSET dataset, the LoRA-enabled ViT-Small model uses the same EO-based obesity fairness objective. Similarly, for the Shenzhen dataset, the same topology minimizes EO disparities between male and female patients.
All systems within the same dataset share identical training procedures and hyperparameters to ensure comparability. For PHN-based systems, optimization is performed using Adam optimizer with a learning rate of 1e41e-4 and batch size of $256$. DenseNet- and LoRA-ViT-Small-based systems employ dataset-specific but fixed hyperparameters across System1 and System2 using Adam optimizer with a batch size of $16$. However, learning rates differ as 5e55e-5 and 1e41e-4 for the mBRSET and Shenzhen datasets, respectively. This shared setup ensures that differences in observed performance arise solely from the fairness objective being optimized to provide a controlled analysis of utility-fairness trade-offs.
4.3 Evaluation Results
(a) Pareto plot.
(b) Radar chart.